# RoFt-Mol: Benchmarking Robust Fine-Tuning with Molecular Graph Foundation Models

### v2

@@ 1 Introduction
In recent years, foundation models (Bommasani et al., 2021; Zhou et al., 2023) have achieved success in learning high-quality, general-purpose representations of images and text through pre-training on diverse datasets (Radford et al., 2021; Kirillov et al., 2023; Ramesh et al., 2022; Touvron et al., 2023; Bubeck et al., 2023; Zhao et al., 2023). To adapt these pre-trained models for downstream applications, additional training on task-specific data, known as fine-tuning, is often required. However, vanilla fine-tuning frequently encounters challenges, including model overfitting (Howard and Ruder, 2018; Li et al., 2020a; Kornblith et al., 2019), catastrophic forgetting of pre-trained knowledge (Lee et al., 2022; Li et al., 2019a; Xuhong et al., 2018; Lubana et al., 2022), and distribution shifts between fine-tuned and test samples, which can lead to negative transfer (Wang et al., 2019; Chen et al., 2019). These challenges highlight the need for robust fine-tuning strategies (Shen et al., 2021; Wortsman et al., 2022; Kumar et al., 2022; Shu et al., 2023; Andreassen et al., 2021; Kirichenko et al., 2022).
Recently, the advantages of foundation models have been extended to various scientific applications (Golling et al., 2024; Leung and Bovy, 2024; Nguyen et al., 2023). Among these, molecular graph foundation models (MGFMs) have gained significant attention for their promising potential in biochemistry (Hu et al., 2020a; Hou et al., 2022a; Xia et al., 2023a; Suresh et al., 2021; Shoghi et al., 2023; Beaini et al., 2023; Zheng et al., 2023; Ross et al., 2022; Rong et al., 2020; Mao et al., 2024). While MGFMs exhibit scaling behaviors similar to foundation models in other domains (Chen et al., 2024), they face unique challenges related to data and tasks.
A primary challenge stems from the significantly smaller pre-training datasets in this domain, typically consisting of at most molecular samples, compared to the billions of samples used in other domains (Sun et al., 2022). This limitation restricts the parameter scale of MGFMs ( parameters) and their generalization capacity (Wang et al., 2024; Akhondzadeh et al., 2023). Furthermore, downstream tasks in this domain often involve limited data for fine-tuning, with datasets containing only tens or a few hundred labeled samples (Wijaya et al., 2024), exacerbating the difficulty of achieving robust model generalization. In addition to data constraints, many downstream tasks, such as molecular property prediction, are regression-based (Wu et al., 2018; Hou et al., 2022b). These tasks require models to capture fine-grained numerical patterns, which presents a distinct requirement compared to the coarse-grained feature reliance typical in classification tasks in CV and NLP. These factors collectively highlight the need for a careful examination of fine-tuning strategies for MGFMs and their appropriate improvement.
@@ 2 Finetuning Methods for Evaluation
Self-supervised Pre-training strategies have been proven to be effective in generating transferable molecular representations for downstream tasks (Zhao et al., 2024). In a high level, they can be divided into reconstruction methods and contrastive methods. The generative-based strategies adopt mask-based graph reconstruction by utilizing graph autoencoders (Hou et al., 2022a; Tan et al., 2023; Wang et al., 2017a; Pan et al., 2018), context predictions (Hu et al., 2020a; Rong et al., 2020) and generative language model pre-training (Hu et al., 2020b; Zhang et al., 2021a). On the other hand, contrastive-based methods aim for maximizing the similarity between perturbed instance pairs (Veličković et al., 2018; Suresh et al., 2021; You et al., 2020; Xia et al., 2023b; Wang et al., 2022; Zhu et al., 2022; You et al., 2021; Qiu et al., 2020; Li et al., 2022; Xu et al., 2021). Moreover, the advancement of language models has prompted numerous studies to employ multi-modal frameworks. These approaches harness language models to enhance molecular understanding through techniques such as cross-modal contrastive learning and alignment (Su et al., 2022; Liu et al., 2023a; Seidl et al., 2023; Liu et al., 2023b).
In this work, we select GraphMAE (Hou et al., 2022a) as the representative of the recontruction-based pre-trained model, which focuses on masked feature reconstruction with scaled cosine error that enabled robust training. Regarding the contrastive pre-trained model, we choose Mole-BERT (Xia et al., 2023b) that combines the node-level masked atom modeling to predict the masked atom tokens and the graph-level contrastive learning through triplet loss and contrastive loss. Lastly, we choose MoleculeSTM (Liu et al., 2023a) as the representative of multi-modal molecule structure-text model that jointly learning molecules’ chemical structures and textual descriptions via a contrastive learning strategy.
Supervised Pre-training. Recently, to leverage more diverse datasets and tasks, researchers started exploring the ability of supervised pre-training with multi-task learning for molecular representations (Gasteiger et al., 2022; Shoghi et al., 2023; Beaini et al., 2023). We adopt pre-trained models trained on multi-task labeled samples in a supervised manner from the Graphium library (Beaini et al., 2023). In addition to the GNN-based backbone, more expressive architectures like Graph Transformer (Ying et al., 2021; Rampášek et al., 2022; Mialon et al., 2021) have been proposed and can be used as the pre-trained backbone with supervised labels, which we adopt GraphGPS Rampášek et al. (2022) as a representative.
@@
Partial FT strategies only optimizes partial weights of the pre-trained model, i.e., a subset of weights within will be updated following the same objective as Eq. 1. Linear Probing (LP) only trains the additional prediction head during the FT. Surgical FT (Lee et al., 2022) updates only partial layers within the encoder. For instance, we can update the weights for -th layer of the GNN encoder as , where is the hyperparameter that can be tuned. LP-FT (Kumar et al., 2022) aims to address the issue of pre-trained feature distortion during the full-FT process. It first performs the LP step to the prediction head while keeping the encoder with fixed pre-trained parameters , followed by applying full-FT with the tuned prediction head.
Weight-based FT strategies mainly update the entire model weights through combining pre-trained model weights and fine-tuned model weights. WiSE-FT (Wortsman et al., 2022) linearly interpolates between pre-training parameters and fine-tuning parameters using a mixing coefficient , to get the interpolated GNN with weights . We first perform full-FT to obtain the adapted encoder and classifier , then apply post-hoc weight ensembling to get , with final predictions given by . , as a hyperparameter, controls the weight ensemble. -SP (Xuhong et al., 2018) regularizes the fine-tuning model weights closer to the pre-trained weights by . We optimize for and by combining the prediction loss from Eq. 1 and with tunable trade-off coefficient .
Representation-based FT methods mainly regulate the latent representation space during FT. Feature-map (Li et al., 2019a) adds distance regularization between the latent representations of pre-trained and fine-tuned models to the Full-FT loss. The regularization is defined as , where controls the regularization strength. BSS (Chen et al., 2019) aims at resolving the negative transfer issue through eliminating the spectral components corresponding to small singular values that are less transferable. The regularization is done as , where is the feature matrix of a batch of graphs and are the -th smallest singular values obtained from the SVD of . We can tune and to determine the number of singular values to penalize and the degree of penalty.
@@ 3 Experimental Settings in the Benchmark
Foundation Models. For self-supervised pre-training, we adopt three open-source pre-trained checkpoints: Mole-BERT, GraphMAE, and MoleculeSTM. For supervised pre-training, we use models from the Graphium (Beaini et al., 2023) library, which get pre-trained on the Toymix and Largemix datasets provided in this library. To differentiate between them, we refer to these models as Graphium-Toy and Graphium-Large. For larger graph transformer based model, we adopt the pre-trained checkpoint of GraphGPS Rampášek et al. (2022) pre-trained on the PCQM4MV2 Hu et al. . For details of datasets used in pre-training are in Appendix D. Furthermore, we include the traditional baseline XGBoost Chen and Guestrin (2016) for Fewshot scenarios to better compare with the foundation model in Appendix G.2.
@@
Dataset Splits. For each downstream dataset, we experiment with random, scaffold, and size splits to create the Train/Val/Test subsets. Specifically, the random splitting shuffles the data, maintaining the Train/Val/Test sets as in-distribution (ID). The other two splitting methods simulate out-of-distribution (OOD) challenges in real-world applications. For scaffold splitting, we follow prior works (Ramsundar et al., 2019), ensuring structural differences in molecular scaffolds across splits. Size splitting, following Zou et al. (2024), arranges molecules in ascending order by size, evaluating model generalization across different molecule sizes.
Size of fine-tuning samples. In practice, molecular property prediction tasks can have very limited experimentally-validated data, e.g., with less than 100 samples (Wijaya et al., 2024). Thus, we consider both Non-Fewshot and Fewshot settings to better simulate the label scarcity issue. In the Non-Fewshot setting, we use all available samples from the splitted train set. In the Fewshot settings, we sample subsets of 50, 100, and 500 molecules from the Train set for fine-tuning, while keeping the Val/Test sets unchanged to ensure a fair comparison. Note that we exclude MUV, Tox21, and ToxCast datasets for the Fewshot settings, as we cannot randomly select training samples while ensuring that all tasks have a specified number of labels simultaneously, due to the severe label scarcity issues in these datasets.
@@ 4.1 Self-supervised Pre-trained Models
When checking the few-shot results in Fig. 2a and 2c, full fine-tuning ranks the highest for regression tasks but only achieves mid-tier performance for classification tasks. This disparity likely arises from the distinct nature of these tasks. Classification tasks typically require coarser-grained features, as exemplified by the Tox21 dataset. In this case, determining toxicity may largely rely on recognizing certain functional groups, such as toxicophores or structural alerts (Singh et al., 2016). In contrast, regression tasks demand finer-grained features. For example, predicting precise solubility involves factors such as partial charge distribution, conformational flexibility, and hydrogen bond patterns, among others (Faller and Ertl, 2007). Consequently, models fine-tuned for regression tasks must acquire more downstream knowledge during the fine-tuning process and are generally less prone to overfitting compared to those used for classification tasks.
@@
As shown in Tables 2 and 3, LP is consistently the worst performing method for self-supervised pre-trained models across all data splits, even under the few-shot fine-tuning in Fig. 2a and 2c. Furthermore, this behavior is widely observed across all tested self-supervised models as GraphMAE and MoleculeSTM, which contrasts the observations in CV where LP demonstrates robust OOD performance by preserving high quality and generalizable features from pre-trained embeddings (Wortsman et al., 2022; Kumar et al., 2022). We attribute this to the misalignment between general-purpose representations produced by self-supervised pre-training and the features required by the specific molecular tasks. Consequently, relying solely on tuning the classifier is insufficient to extract meaningful predictions from these non-informative representations.
@@
Among various fine-tuning methods, weight-based approaches consistently outperform others across a wide range of experiments, regardless of the few-shot sample sizes (Fig. 2a and 2c). Self-supervised models are known to capture general-purpose knowledge for substructure discovery(Wang et al., 2024). During fine-tuning, combining pre-trained and fine-tuned weights proves effective in extracting molecular patterns relevant to downstream tasks. Notably, WiSE-FT demonstrates superior performance on classification datasets, whereas -SP excels in regression tasks. This finding is also supported by MoleculeSTM in table 11 where -SP remains as top method under all few-shot regression tasks and WiSE-FT excels under Fewshot-50 classification. Essentially, WiSE-FT applies a straightforward post-hoc linear interpolation between pre-trained and fine-tuned models, governed by a single coefficient. In contrast, -SP implicitly determines the weight combination through the training loss (Lubana et al., 2022; Xuhong et al., 2018), aligning with statement (1a) that regression tasks typically demand more nuanced modeling.
@@ 5.2 Algorithm
This approach naturally incorporates the characteristics of -SP and even surgical FT: The weight ensemble in DWiSE-FT offers the flexibility through varying mixing layer-wise coefficients between the pre-trained and fine-tuned models, addressing the limitations of WiSE-FT. Additionally, we enable the selection of through optimization via validation loss gradient inspired by the Gradient-based Neural Architecture Search (NAS) (Dong and Yang, 2019).

### v3

@@ 1 Introduction
In recent years, foundation models [bommasani2021opportunities, zhou2023comprehensive] have achieved success in learning high-quality, general-purpose representations of images and text through pre-training on diverse datasets [radford2021learning, kirillov2023segment, ramesh2022hierarchical, touvron2023llama, bubeck2023sparks, zhao2023survey]. To adapt these pre-trained models for downstream applications, additional training on task-specific data, known as fine-tuning, is often required. However, vanilla fine-tuning frequently encounters challenges, including model overfitting [howard2018universal, lirethinking, kornblith2019better], catastrophic forgetting of pre-trained knowledge [leesurgical, lidelta, xuhong2018explicit, lubana2022quadratic], and distribution shifts between fine-tuned and test samples, which can lead to negative transfer [wang2019characterizing, chen2019catastrophic]. These challenges highlight the need for robust fine-tuning strategies [shen2021partial, wortsman2022robust, kumar2022fine, shu2023clipood, andreassenevolution, kirichenkolast].
Recently, the advantages of foundation models have been extended to various scientific applications [golling2024masked, leung2024towards, nguyen2023climax]. Among these, molecular graph foundation models (MGFMs) have gained significant attention for their promising potential in biochemistry [hu2020strategies, hou2022graphmae, xiamole, suresh2021adversarial, shoghimolecules, beainitowards, zheng2023towards, ross2022large, rong2020self, maoposition]. While MGFMs exhibit scaling behaviors similar to foundation models in other domains [chen2024uncovering], they face unique challenges related to data and tasks.
A primary challenge stems from the significantly smaller pre-training datasets in this domain, typically consisting of at most molecular samples, compared to the billions of samples used in other domains [sun2022does]. This limitation restricts the parameter scale of MGFMs ( parameters) and their generalization capacity [wang2024evaluating, akhondzadeh2023probing]. Furthermore, downstream tasks in this domain often involve limited data for fine-tuning, with datasets containing only tens or a few hundred labeled samples [wijaya2024two], exacerbating the difficulty of achieving robust model generalization. In addition to data constraints, many downstream tasks, such as molecular property prediction, are regression-based [wu2018moleculenet, hou2022accurate]. These tasks require models to capture fine-grained numerical patterns, which presents a distinct requirement compared to the coarse-grained feature reliance typical in classification tasks in CV and NLP. These factors collectively highlight the need for a careful examination of fine-tuning strategies for MGFMs and their appropriate improvement.
@@ 2 Finetuning Methods for Evaluation
Self-supervised Pre-training strategies have been proven to be effective in generating transferable molecular representations for downstream tasks [zhao2024survey]. In a high level, they can be divided into reconstruction methods and contrastive methods. The generative-based strategies adopt mask-based graph reconstruction by utilizing graph autoencoders [hou2022graphmae, tan2023s2gae, wang2017mgae, pan2018adversarially], context predictions [hu2020strategies, rong2020self] and generative language model pre-training [hu2020gpt, zhang2021motif]. On the other hand, contrastive-based methods aim for maximizing the similarity between perturbed instance pairs [velivckovic2018deep, suresh2021adversarial, you2020graph, xia2023mole, wang2022molecular, zhu2022unified, you2021graph, qiu2020gcc, li2022let, xu2021self]. Moreover, the advancement of language models has prompted numerous studies to employ multi-modal frameworks. These approaches harness language models to enhance molecular understanding through techniques such as cross-modal contrastive learning and alignment [su2022molecular, liu2023multi, seidl2023enhancing, liu2023molca].
In this work, we select GraphMAE [hou2022graphmae] as the representative of the recontruction-based pre-trained model, which focuses on masked feature reconstruction with scaled cosine error that enabled robust training. Regarding the contrastive pre-trained model, we choose Mole-BERT [xia2023mole] that combines the node-level masked atom modeling to predict the masked atom tokens and the graph-level contrastive learning through triplet loss and contrastive loss. Lastly, we choose MoleculeSTM [liu2023multi] as the representative of multi-modal molecule structure-text model that jointly learning molecules’ chemical structures and textual descriptions via a contrastive learning strategy.
Supervised Pre-training. Recently, to leverage more diverse datasets and tasks, researchers started exploring the ability of supervised pre-training with multi-task learning for molecular representations [gasteigergemnet, shoghimolecules, beainitowards]. We adopt pre-trained models trained on multi-task labeled samples in a supervised manner from the Graphium library [beainitowards]. In addition to the GNN-based backbone, more expressive architectures like Graph Transformer [ying2021transformers, rampavsek2022recipe, mialon2021graphit] have been proposed and can be used as the pre-trained backbone with supervised labels, which we adopt GraphGPS [rampavsek2022recipe] as a representative.
@@
Partial FT strategies only optimizes partial weights of the pre-trained model, i.e., a subset of weights within will be updated following the same objective as Eq. 1. Linear Probing (LP) only trains the additional prediction head during the FT. Surgical FT [leesurgical] updates only partial layers within the encoder. For instance, we can update the weights for -th layer of the GNN encoder as , where is the hyperparameter that can be tuned. LP-FT [kumar2022fine] aims to address the issue of pre-trained feature distortion during the full-FT process. It first performs the LP step to the prediction head while keeping the encoder with fixed pre-trained parameters , followed by applying full-FT with the tuned prediction head.
Weight-based FT strategies mainly update the entire model weights through combining pre-trained model weights and fine-tuned model weights. WiSE-FT [wortsman2022robust] linearly interpolates between pre-training parameters and fine-tuning parameters using a mixing coefficient , to get the interpolated GNN with weights . We first perform full-FT to obtain the adapted encoder and classifier , then apply post-hoc weight ensembling to get , with final predictions given by . , as a hyperparameter, controls the weight ensemble. -SP [xuhong2018explicit] regularizes the fine-tuning model weights closer to the pre-trained weights by . We optimize for and by combining the prediction loss from Eq. 1 and with tunable trade-off coefficient .
Representation-based FT methods mainly regulate the latent representation space during FT. Feature-map [lidelta] adds distance regularization between the latent representations of pre-trained and fine-tuned models to the Full-FT loss. The regularization is defined as , where controls the regularization strength. BSS [chen2019catastrophic] aims at resolving the negative transfer issue through eliminating the spectral components corresponding to small singular values that are less transferable. The regularization is done as , where is the feature matrix of a batch of graphs and are the -th smallest singular values obtained from the SVD of . We can tune and to determine the number of singular values to penalize and the degree of penalty.
@@ 3 Experimental Settings in the Benchmark
Foundation Models. For self-supervised pre-training, we adopt three open-source pre-trained checkpoints: Mole-BERT, GraphMAE, and MoleculeSTM. For supervised pre-training, we use models from the Graphium [beainitowards] library, which get pre-trained on the Toymix and Largemix datasets provided in this library. To differentiate between them, we refer to these models as Graphium-Toy and Graphium-Large. For larger graph transformer based model, we adopt the pre-trained checkpoint of GraphGPS [rampavsek2022recipe] pre-trained on the PCQM4MV2 [hu2ogb]. For details of datasets used in pre-training are in Appendix D. Furthermore, we include the traditional baseline XGBoost [chen2016xgboost] for Fewshot scenarios to better compare with the foundation model in Appendix G.2.
@@
Dataset Splits. For each downstream dataset, we experiment with random, scaffold, and size splits to create the Train/Val/Test subsets. Specifically, the random splitting shuffles the data, maintaining the Train/Val/Test sets as in-distribution (ID). The other two splitting methods simulate out-of-distribution (OOD) challenges in real-world applications. For scaffold splitting, we follow prior works [ramsundar2019deep], ensuring structural differences in molecular scaffolds across splits. Size splitting, following zougess, arranges molecules in ascending order by size, evaluating model generalization across different molecule sizes.
Size of fine-tuning samples. In practice, molecular property prediction tasks can have very limited experimentally-validated data, e.g., with less than 100 samples [wijaya2024two]. Thus, we consider both Non-Fewshot and Fewshot settings to better simulate the label scarcity issue. In the Non-Fewshot setting, we use all available samples from the splitted train set. In the Fewshot settings, we sample subsets of 50, 100, and 500 molecules from the Train set for fine-tuning, while keeping the Val/Test sets unchanged to ensure a fair comparison. Note that we exclude MUV, Tox21, and ToxCast datasets for the Fewshot settings, as we cannot randomly select training samples while ensuring that all tasks have a specified number of labels simultaneously, due to the severe label scarcity issues in these datasets.
@@ 4.1 Self-supervised Pre-trained Models
When checking the few-shot results in Fig. 2a and 2c, full fine-tuning ranks the highest for regression tasks but only achieves mid-tier performance for classification tasks. This disparity likely arises from the distinct nature of these tasks. Classification tasks typically require coarser-grained features, as exemplified by the Tox21 dataset. In this case, determining toxicity may largely rely on recognizing certain functional groups, such as toxicophores or structural alerts [singh2016toxicophore]. In contrast, regression tasks demand finer-grained features. For example, predicting precise solubility involves factors such as partial charge distribution, conformational flexibility, and hydrogen bond patterns, among others [faller2007computational]. Consequently, models fine-tuned for regression tasks must acquire more downstream knowledge during the fine-tuning process and are generally less prone to overfitting compared to those used for classification tasks.
@@
As shown in Tables 2 and 3, LP is consistently the worst performing method for self-supervised pre-trained models across all data splits, even under the few-shot fine-tuning in Fig. 2a and 2c. Furthermore, this behavior is widely observed across all tested self-supervised models as GraphMAE and MoleculeSTM, which contrasts the observations in CV where LP demonstrates robust OOD performance by preserving high quality and generalizable features from pre-trained embeddings [wortsman2022robust, kumar2022fine]. We attribute this to the misalignment between general-purpose representations produced by self-supervised pre-training and the features required by the specific molecular tasks. Consequently, relying solely on tuning the classifier is insufficient to extract meaningful predictions from these non-informative representations.
@@
Among various fine-tuning methods, weight-based approaches consistently outperform others across a wide range of experiments, regardless of the few-shot sample sizes (Fig. 2a and 2c). Self-supervised models are known to capture general-purpose knowledge for substructure discovery[wang2024evaluating]. During fine-tuning, combining pre-trained and fine-tuned weights proves effective in extracting molecular patterns relevant to downstream tasks. Notably, WiSE-FT demonstrates superior performance on classification datasets, whereas -SP excels in regression tasks. This finding is also supported by MoleculeSTM in table 11 where -SP remains as top method under all few-shot regression tasks and WiSE-FT excels under Fewshot-50 classification. Essentially, WiSE-FT applies a straightforward post-hoc linear interpolation between pre-trained and fine-tuned models, governed by a single coefficient. In contrast, -SP implicitly determines the weight combination through the training loss [lubana2022quadratic, xuhong2018explicit], aligning with statement (1a) that regression tasks typically demand more nuanced modeling.
@@ 5.2 Algorithm
This approach naturally incorporates the characteristics of -SP and even surgical FT: The weight ensemble in DWiSE-FT offers the flexibility through varying mixing layer-wise coefficients between the pre-trained and fine-tuned models, addressing the limitations of WiSE-FT. Additionally, we enable the selection of through optimization via validation loss gradient inspired by the Gradient-based Neural Architecture Search (NAS) [dong2019searching].